OpenStudy (anonymous):

suppose an investment of 7,300 dubles in value ever nine years. how much is the investment worth after 27 years

OpenStudy (texaschic101):

9 years = (7300)(2) = 14,600 18 years = (14,600)(2) = 29,200 27 years = (29,200)(2) = 58,400 I guess I did this correctly. You might want to wait for a second opinion.

OpenStudy (whpalmer4):

Doubling every period P is given by \[2^{t/P}\]where \(t\) and \(P\) are in identical units. If our investment doubles every 9 years, and starts out at 7300, the amount after \(t\) years is \[7300*2^{t/9}\]and after 27 years yields\[7300*2^{27/9} = 7300*2^3=7300*8=58400\]

OpenStudy (whpalmer4):

That works out to an annual compounded growth rate of just over 8%.